A train staring from a railway station  and moving with uniform acceleration ,attains a speed of 40km/h in 10 minutes.its acceleration  is

Grade:9

1 Answers

geet kasyap
6 years ago
the train starts from rest
u=0 
And gains speed to 40
v=40km/h
T=10min=1/6hr
 
Acceleration=v-u/t
Acc=40×6. (Direct step)
Acceleration =240kmph